Find the domain:
g(x)= (3x+15)/(x^2+8x-20)

Answer by oscargut(2103)

Domain = x/ (x^2+8x-20 is not 0)
x^2+8x-20 = 0
(x+10)(x-2) = 0
x = -10 or x = 2
Answer: Domain = {x/ x is not 2, x is not -10}
If you have more questions i can see them if you contact me at: mthman@gmail.com
Thanks